OpenStudy (anonymous):

Find all points on the circle of radius 6 centered at the origin where the slope is ¾. Show all steps.

OpenStudy (anonymous):

A circle does not have a slope. Do you need the intersection of the circle and a line with slope 3/4?

OpenStudy (anonymous):

Well more accurately find the points at which the tangent lines have a slope of 3/4.

OpenStudy (anonymous):

the equation of the circle will be: \[x^2+y^2=36\]

OpenStudy (anonymous):

now expressing it as a function: \[y=\sqrt{(36-x^2)}\] and \[y=-\sqrt{(36-x^2)}\]

OpenStudy (anonymous):

now you take the derivative of the functions: \[y'=-2x(36-x^2)^(-1/2)\] and \[y'=2x(36-x^2)^(-1/2)\]

OpenStudy (anonymous):

or you can take the derivative directly and get \[2x+2yy'=0\] so \[y'=-\frac{x}{y}\]

OpenStudy (anonymous):

clever!

OpenStudy (anonymous):

@satellite73 i got that but how u get points from there?

OpenStudy (anonymous):

not done yet though (thanks)

OpenStudy (anonymous):

with my technique you just have to solve it when y'=3/4

OpenStudy (anonymous):

ya but im supposed to do the prob the calc way...but ty anyway!!! calc bc is a pain...

OpenStudy (anonymous):

both ways is calculus

OpenStudy (anonymous):

ok well using implicit diff like satellite is doing

OpenStudy (anonymous):

that means if \[(x,y)\] is on the circle then \[-\frac{x}{y}=\frac{3}{4}\] or \[-4x=3y\] or \[y=-\frac{4}{3}x\] so write \[x^2+(-\frac{4}{3}x)^2=36\] and solve for x

OpenStudy (anonymous):

wow that is nice. thanks